Jack stride is 64 cm long & Jills is 56 cm long. If they dtart together,how far have they walked when they are in next step?

Answer by dfrazzetto(283)

ratio of strides:
64/56 = 8*8/8*7 = 8/7
To bring back into "step" 8/7 * 7/8 = 1
Jill has to walk 8 strides for every 7 Jack walks so:
56 * 8 = 64 * 7 = 448 cm